# Grounding Foundation Models through Federated Transfer Learning: A General Framework

### v10

@@ 4. FTL-FM Taxonomy
From the perspective of data and model protection, most of the FTL-FM works (superscripted by ***) focus on protecting data privacy under semi-honest attacks. Backdoor (superscripted by $\ddagger$), including both attack and defense, is also an active research area. While the model ownership protection (superscripted by $\dagger$) is under-explored.
@@
Semi-Honest $\mathcal{D}^{S}$ ($\mathcal{W}^{S}$) PPDG  (Carranza et al., 2023; Kurakin et al., $*{}^{*}$  (Mattern et al., 2022; Yue et al., $*{}^{*}$ DP  (Behnia et al., 2022; Bu et al., 2022; Shi et al., $*{}^{*}$, MP (Youssef et al., $*{}^{*}$ DP  (Yu et al., 2022; Mireshghallah et al., $*{}^{*}$ CP (Xiao et al., 2023)  (Fan et al., 2023; Chua et al., 2023; Khalid et al., ${}^{\dagger}$
$\mathcal{D}^{C}$ ($\mathcal{W}^{C}$) DP (Li et al., 2023c; Shen et al., $*{}^{*}$, WPE (Xu et al., $*{}^{*}$, PTB (Zhou et al., $*{}^{*}$ PPDS (Hou et al., 2023b; Yu et al., $*{}^{*}$ DP (Wang et al., $*{}^{*}$ SA (Fan et al., $*{}^{*}$ DP (Azam et al., 2023; Balunovic et al., $*{}^{*}$, FE (Gupta et al., $*{}^{*}$ HE (Jin et al., $*{}^{*}$ DP  (Xu et al., 2023a; Zhao et al., $*{}^{*}$, PEFT (Zhang et al., $*{}^{*}$, SA (Fan et al., $*{}^{*}$ TEE (Huang et al., ${}^{*\dagger}$
Malicious $\mathcal{D}^{S}$ ($\mathcal{W}^{S}$) NC, DP (Zhang et al., 2022b; Yoo and Kwak, ${}^{\ddagger}$
$\mathcal{D}^{C}$ ($\mathcal{W}^{C}$) FP (Zhang et al., 2021; Shen et al., ${}^{\ddagger}$, OWD (Chen et al., 2022b; Mei et al., [n. ${}^{\ddagger}$, MDP (Xi et al., ${}^{\ddagger}$ OWD (Cai et al., ${}^{\ddagger}$ DP (Chu et al., $*{}^{*}$  (Fowl et al., 2023; Rashid et al., $*{}^{*}$ SCR (Rashid et al., $*{}^{*}$
@@
*** signifies works defending against data privacy attacks. $\dagger$ signifies works defending against model stealing attacks. $\ddagger$ signifies works defending against backdoor attacks.
@@ 5. Foundation Models for Enhancing Domain Models
In this section, we overview FTL-FM methods that fall into setting 1⃝ (see Definition 3.1), the objective of which is to transfer and adapt the knowledge of foundation models (FMs) to downstream clients’ domain models (DMs) for exploiting the power of FMs. To this end, the literature proposed a variety of methods involving data-level knowledge transfer, representation-level knowledge transfer, and model-level knowledge transfer.
@@
Fr. $S_{\text{FM}}$:
$\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$
Fr. $S_{\text{Fed}}$:
Global model
To $S_{\text{FM}}$:
$X$ + soft $P$
To $S_{\text{Fed}}$:
$\mathcal{W}^{C}$
@@
Fr. $S_{\text{FM}}$:
$\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$
Fr. $S_{\text{Fed}}$:
Global distr. of $P$
To $S_{\text{FM}}$:
$X$ + soft $P$
To $S_{\text{Fed}}$:
Local distr. of $P$
@@
Fr. $S_{\text{FM}}$:
$\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$
Fr. $S_{\text{Fed}}$:
Global encoder
To $S_{\text{FM}}$:
Perturbed $X$
To $S_{\text{Fed}}$:
Local encoder
@@ 5.1. FM adapted to DMs through data-level knowledge transfer
The literature has explored mainly four categories of data-level knowledge transfer methods that adapt the knowledge of FMs to a client of a specific domain: (1) the client prompts tuning a closed-source FM (e.g., ChatGPT) by optimizing a local prompt generator based on its domain-specific data and predictions generated by the FM. (2) the client optimizes discrete prompts with the guidance of FMs. (3) the client fine-tunes its domain model with knowledge generated by FMs. (4) An FM server transfers data generated by general-purpose FMs or FMs augmented with industry-level knowledge (e.g., LawGPT (Nguyen, 2023), FinGPT (Yang et al., 2023), BloombergGPT (Wu et al., 2023a), and Med-PaLM2 (Singhal et al., 2023)) to clients of specific domains. We summarize data-level knowledge transfer methods that fall into setting 1⃝ in Table 3222While we consider an FTL-FM approach should incorporate certain protection mechanisms to protect data or models, we still review highly relevant works that prioritize the development of knowledge transfer methods without considering data or model protection and include them in the Table to ensure thoroughness. (same for Section 6 and Section 7). These methods leave ”blank” in the ”What To Protect” and ”How To Protect” columns. and elaborate on these methods in this subsection.
@@ 5.1.1. Prompt Generator Optimization
Alternatively, clients can adapt knowledge of the FM to their specific domains through federated prompt optimization. A typical setting of federated prompt optimization involves a client that learns a local prompt generator to generate prompts that can guide the FM toward the desired responses (illustrated in Figure 6(a)). For example, BBT (Sun et al., 2022b), BBTv2 (Sun et al., 2022a), and BlackVIP (Oh et al., 2023) proposed that a client can customize a remote FM (e.g., ${\text{RoBERTa}}_{\text{large}}$ (Liu et al., 2019), ${\text{BART}}_{\text{large}}$ (Lewis et al., 2020), and ViT-B (Radford et al., 2021)) by optimizing a domain-specific prompt generator based on its local data together with the predictions (e.g., logits) generated by the remote FM. BBT and BBTv2 employ a gradient-free method called CMA-ES (Covariance Matrix Adaptation Evolution Strategy) (Hansen, 2016) to optimize the client’s local prompt generator, whereas BlackVIP leverages SPSA (Simultaneous Perturbation Stochastic Approximation) (Spall, 1992) to efficiently approximates the high-dimensional gradients, which then are used to optimize the prompt generator. BBT, BBTv2, and BlackVIP require the FM API to take continuous (i.e., soft) prompts as input, and thus, they cannot be applied to FM APIs that only accept discrete inputs. To eliminate this constraint, BDPL (Diao et al., 2023) and DSP (Li et al., 2023b) were proposed to learn a client side’s discrete (i.e., hard) prompt generator (DPG) for customizing FM’s responses. BDPL adopts a variance-reduced policy gradient algorithm (Zhou et al., 2021) to optimize the DPG, while DSP optimizes the DPG through supervised fine-tuning and reinforcement learning. BBT, BBTv2, BlackVIP, BDPL, and DSP involve only one client, and thus, they correspond to stage 2 of the first objective formulated in Eq.(2) with $K=1,p^{C_{1}}=1$, $\eta=0$, the domain model $\mathcal{W}^{C_{1}}$ is the client’s local prompt generator and the specific form of the loss $\ell^{C_{1}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{1}};d)$ is formulated in Eq.(7).
@@
The aforementioned federated prompt tuning methods involve an FM server and only one client. Fed-BBPT (Lin et al., 2023) and FedBPT (Sun et al., 2023) extend the federated prompt tuning to multiple clients, the workflow of which is illustrated in Figure  6(b). More specifically, Fed-BBPT and FedBPT involve two servers: an FM server, denoted as $S_{\text{FM}}$, and a federated server, denoted as $S_{\text{Fed}}$. The former hosts an FM to be prompted by clients, while the latter is for federated tuning clients’ local prompt generators. In Fed-BBPT, each client leverages the Simultaneous Perturbation Stochastic Approximation (Spall, 1997) to train its prompt generators based on local data and predictions generated by the FM, and the federated server aggregates clients’ local prompt generators to form the global prompt generator. FedBPT differs from Fed-BBPT in that each client in FedBPT exploits CMA-ES to optimize the distribution of prompts generated by its prompt generator, and the server derives the global prompt generator by aggregating local distributions of clients’ prompts. Generally, Fed-BBPT and FedBPT adhere to the objective formulated in Eq.(2) with $K>1$, $\eta=0$, the domain model $\mathcal{W}^{C_{k}}$ of each client $k$ is a prompt generator and the specific loss of federated prompt tuning $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ between the FM server and a client $k$ is formulated in Eq.(7).
@@ 5.1.4. Synthetic Data Transfer
In some proprietary domains, such as the audit department in a bank, data resources are typically limited for deep learning tasks. One potential solution to address data scarcity is leveraging FMs pre-trained on a large corpus of domain-specific data to generate synthetic training and testing data (e.g., LawGPT (Nguyen, 2023)). However, domain-specific data (e.g., financial and legal data) typically contain sensitive and private information. Thus, directly releasing FMs trained on these data presents potential privacy issues because sensitive information encoded in FMs can possibly be recovered by adversaries. In order to share synthetic data generated by FMs without compromising the privacy of training data, privacy-preserving data generation methods have been proposed (Carranza et al., 2023; Yue et al., 2023; Mattern et al., 2022; Kurakin et al., 2023). Essentially, these methods first fine-tune a pretrained public FM using differential privacy (DP) on private domain data and then exploit the DP-tuned FM to generate new data, which can be distributed to downstream clients to augment their training data (illustrated in Figure 7(c)). These methods correspond to stage 1 of objective 1⃝ formulated in Eq.(1) but with an additional step of generating data using fine-tuned FMs.
@@ 5.2. FM adapted to DMs through representation-level knowledge transfer
Representation-level knowledge transfer (RLKT) refers to transferring and adapting knowledge of FMs to domain-specific models through passing representation-level knowledge, such as intermediate forward outputs and backward gradients, between FL participating parties. In literature, split learning (Liu et al., 2022) and representation transfer are two main categories of RLKT explored in the setting 1⃝ of FTL-FM. We summarize the two categories of RLKT methods in Table 4 and elaborate on these methods in this subsection.
@@
SAP (Shen et al., 2023) takes a similar setting as TextObfuscator, but it adopts a different training strategy. During training, the client model is frozen, while the server model is trained using LoRA. To protect its data privacy, the client applies $d_{\chi}$-privacy (Chatzikokolakis et al., 2013) to obfuscate representation vectors before transmitting them to the server. Furthermore, SAP studied the impact of the FM’s split position on the privacy-preserving capacity and the model performance of the SAP framework. It empirically demonstrated that as the number of layers in the bottom model increases, the utility diminishes while the privacy-preserving capability strengthens, highlighting the existence of a trade-off between utility and privacy when partitioning the FM. SAP is a variant of the typical federated split learning we formulated in Eq.(8) as it conducts the FM adaptation through fine-tuning the model on the server side, as opposed to the client model.
@@ 5.4. Privacy and Backdoor Attacks
In the setting 1⃝, the server is the potential adversary, who can launch privacy attacks with the intention of inferring clients’ private information. The server can also carry out backdoor attacks to compromise the utility of clients’ domain models.
@@ 6. Domain Knowledge for Augmenting Foundation Models
In this section, we overview FTL-FM methods that fall into setting 2⃝ (see Definition 3.1), the objective of which is to augment the FM with industry-level and domain-specific knowledge (e.g., healthcare, medicine, finance, banking, and law) transferred from clients.
@@
Hetero- Model Transfer Offsite-Tuning (Xiao et al., 2023) $\bigcirc{}^{\dagger}$ Adapter Emulator of $\mathcal{W}^{S}_{\text{FM}}$ CP
FedOST (WeBank, [n. d.]; Fan et al., 2023) $\bigcirc \bigcirc{}^{\dagger}$ Adapter Emulator of $\mathcal{W}^{S}_{\text{FM}}$, Initial adapter SA CP
FedOT (Kuang et al., 2023a) $\bigcirc{}^{\dagger}$ Adapter Emulator of $\mathcal{W}^{S}_{\text{FM}}$, Initial adapter CP
FedPEAT (Chua et al., 2023) $\bigcirc{}^{\dagger}$ Adapter Emulator of $\mathcal{W}^{S}_{\text{FM}}$, Global adapter CP
@@ 6.2. Domain knowledge transferred to FM through model-level knowledge transfer
To alleviate these limitations, Fan et al. (Fan et al., 2023) proposed a federated learning version of Offsite-Tuning named FedOST. In FedOST, the server first distributes the emulator and two adapters of its FM to all clients at the beginning of FL. Then, each client fine-tunes the two adapters on its private data with the assistance of the emulator. Next, all clients send their adapters to the server, which aggregates received adapters through secure aggregation (Bonawitz et al., 2016) and plugs the aggregated adapter into the FM. FedOST protects the privacy of clients’ private data and can build a fine-tuned FM with enhanced generalization capability. With similar motivations as FedOST, Kuang et al. integrated Offsite-Tuning into FederatedScope-LLM (Kuang et al., 2023b) and termed it FedOT. FedPEAT (Chua et al., 2023) generalized the Offsite-Tuning approach to federated learning with multiple communication rounds and proposed an adaptive control mechanism to facilitate the adoption of FedPEAT in a dynamic environment. CEFHRI (Khalid et al., 2023) adopted Offsite-Tuning to the human-robot interaction domain.

### v11

@@ 4. FTL-FM Taxonomy
From the perspective of data and model protection, most of the FTL-FM works (superscripted by *) focus on protecting data privacy under semi-honest attacks. Backdoor (superscripted by $\ddagger$), including both attack and defense, is also an active research area. While the model ownership protection (superscripted by $\dagger$) is under-explored.
@@
Semi-Honest $\mathcal{D}^{S}$ ($\mathcal{W}^{S}$) PPDG  (Carranza et al., 2023; Kurakin et al., 2023)  (Mattern et al., 2022; Yue et al., 2023) DP  (Behnia et al., 2022; Bu et al., 2022; Shi et al., 2022b), MP (Youssef et al., 2023) DP  (Yu et al., 2022; Mireshghallah et al., 2022) CP (Xiao et al., 2023)  (Fan et al., 2023; Chua et al., 2023; Khalid et al., 2023)
$\mathcal{D}^{C}$ ($\mathcal{W}^{C}$) DP (Li et al., 2023c; Shen et al., 2023), WPE (Xu et al., 2023c), PTB (Zhou et al., 2023a) PPDS (Hou et al., 2023b; Yu et al., 2023a) DP (Wang et al., 2023c) SA (Fan et al., 2023; WeBank, [n. d.]) DP (Azam et al., 2023; Balunovic et al., 2022), FE (Gupta et al., 2022) HE (Jin et al., 2023) DP  (Xu et al., 2023a; Zhao et al., 2023b), PEFT (Zhang et al., 2023d), SA (Fan et al., 2023) TEE (Huang et al., 2024)∗†
Malicious $\mathcal{D}^{S}$ ($\mathcal{W}^{S}$) NC, DP (Zhang et al., 2022b; Yoo and Kwak, 2022)
$\mathcal{D}^{C}$ ($\mathcal{W}^{C}$) FP (Zhang et al., 2021; Shen et al., 2021), OWD (Chen et al., 2022b; Mei et al., [n. d.]), MDP (Xi et al., 2023) OWD (Cai et al., 2022) DP (Chu et al., 2023)  (Fowl et al., 2023; Rashid et al., 2023) SCR (Rashid et al., 2023)
@@
* signifies works defending against data privacy attacks. $\dagger$ signifies works defending against model stealing attacks. $\ddagger$ signifies works defending against backdoor attacks.
@@ 5. Foundation Models for Enhancing Domain Models
In this section, we overview FTL-FM methods that fall into setting ① (see Definition 3.1), the objective of which is to transfer and adapt the knowledge of foundation models (FMs) to downstream clients’ domain models (DMs) for exploiting the power of FMs. To this end, the literature proposed a variety of methods involving data-level knowledge transfer, representation-level knowledge transfer, and model-level knowledge transfer.
@@
Fr. $S_{\text{FM}}$: $\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$
Fr. $S_{\text{Fed}}$: Global model
To $S_{\text{FM}}$: $X$ + soft $P$
To $S_{\text{Fed}}$: $\mathcal{W}^{C}$
@@
Fr. $S_{\text{FM}}$: $\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$
Fr. $S_{\text{Fed}}$: Global distr. of $P$
To $S_{\text{FM}}$: $X$ + soft $P$
To $S_{\text{Fed}}$: Local distr. of $P$
@@
Fr. $S_{\text{FM}}$: $\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$
Fr. $S_{\text{Fed}}$: Global encoder
To $S_{\text{FM}}$: Perturbed $X$
To $S_{\text{Fed}}$: Local encoder
@@ 5.1. FM adapted to DMs through data-level knowledge transfer
The literature has explored mainly four categories of data-level knowledge transfer methods that adapt the knowledge of FMs to a client of a specific domain: (1) the client prompts tuning a closed-source FM (e.g., ChatGPT) by optimizing a local prompt generator based on its domain-specific data and predictions generated by the FM. (2) the client optimizes discrete prompts with the guidance of FMs. (3) the client fine-tunes its domain model with knowledge generated by FMs. (4) An FM server transfers data generated by general-purpose FMs or FMs augmented with industry-level knowledge (e.g., LawGPT (Nguyen, 2023), FinGPT (Yang et al., 2023), BloombergGPT (Wu et al., 2023a), and Med-PaLM2 (Singhal et al., 2023)) to clients of specific domains. We summarize data-level knowledge transfer methods that fall into setting ① in Table 3222While we consider an FTL-FM approach should incorporate certain protection mechanisms to protect data or models, we still review highly relevant works that prioritize the development of knowledge transfer methods without considering data or model protection and include them in the Table to ensure thoroughness. (same for Section 6 and Section 7). These methods leave ”blank” in the ”What To Protect” and ”How To Protect” columns. and elaborate on these methods in this subsection.
@@ 5.1.1. Prompt Generator Optimization
Alternatively, clients can adapt knowledge of the FM to their specific domains through federated prompt optimization. A typical setting of federated prompt optimization involves a client that learns a local prompt generator to generate prompts that can guide the FM toward the desired responses (illustrated in Figure 6(a)).
@@
BBT (Sun et al., 2022b), BBTv2 (Sun et al., 2022a), and BlackVIP (Oh et al., 2023) proposed that a client can customize a remote FM (e.g., ${\text{RoBERTa}}_{\text{large}}$ (Liu et al., 2019), ${\text{BART}}_{\text{large}}$ (Lewis et al., 2020), and ViT-B (Radford et al., 2021)) by optimizing a domain-specific prompt generator based on its local data together with the predictions (e.g., logits) generated by the remote FM. BBT and BBTv2 employ a gradient-free method called CMA-ES (Covariance Matrix Adaptation Evolution Strategy) (Hansen, 2016) to optimize the client’s local prompt generator, whereas BlackVIP leverages SPSA (Simultaneous Perturbation Stochastic Approximation) (Spall, 1992) to efficiently approximates the high-dimensional gradients, which then are used to optimize the prompt generator. BBT, BBTv2, and BlackVIP require the FM API to take continuous (i.e., soft) prompts as input, and thus, they cannot be applied to FM APIs that only accept discrete inputs. To eliminate this constraint, BDPL (Diao et al., 2023) and DSP (Li et al., 2023b) were proposed to learn a client side’s discrete (i.e., hard) prompt generator (DPG) for customizing FM’s responses. BDPL adopts a variance-reduced policy gradient algorithm (Zhou et al., 2021) to optimize the DPG, while DSP optimizes the DPG through supervised fine-tuning and reinforcement learning. BBT, BBTv2, BlackVIP, BDPL, and DSP involve only one client, and thus, they correspond to stage 2 of the first objective formulated in Eq.(2) with $K=1,p^{C_{1}}=1$, $\eta=0$, the domain model $\mathcal{W}^{C_{1}}$ is the client’s local prompt generator and the specific form of the loss $\ell^{C_{1}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{1}};d)$ is formulated in Eq.(8).
@@
The aforementioned federated prompt tuning methods involve an FM server and only one client. Fed-BBPT (Lin et al., 2023) and FedBPT (Sun et al., 2023) extend the federated prompt tuning to multiple clients, the workflow of which is illustrated in Figure  6(b). More specifically, Fed-BBPT and FedBPT involve two servers: an FM server, denoted as $S_{\text{FM}}$, and a federated server, denoted as $S_{\text{Fed}}$. $S_{\text{FM}}$ hosts an FM to be prompted by clients, whereas $S_{\text{Fed}}$ is for federated tuning clients’ local prompt generators. In Fed-BBPT, each client leverages the Simultaneous Perturbation Stochastic Approximation (Spall, 1997) to train its prompt generators based on local data and predictions generated by the FM, and the federated server aggregates clients’ local prompt generators to form the global prompt generator. FedBPT differs from Fed-BBPT in that each client in FedBPT exploits CMA-ES to optimize the distribution of prompts generated by its prompt generator, and the server derives the global prompt generator by aggregating local distributions of clients’ prompts. Generally, Fed-BBPT and FedBPT adhere to the objective formulated in Eq.(2) with $K>1$, $\eta=0$, the domain model $\mathcal{W}^{C_{k}}$ of each client $k$ is a prompt generator and the specific loss of federated prompt tuning $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ between the FM server and a client $k$ is formulated in Eq.(8).
@@ 5.1.4. Synthetic Data Transfer
In some proprietary domains, such as the audit department in a bank, data resources are typically limited for deep learning tasks. One potential solution to address data scarcity is leveraging FMs pre-trained on a large corpus of domain-specific data to generate synthetic training and testing data (e.g., LawGPT (Nguyen, 2023)). However, domain-specific data (e.g., financial and legal data) typically contain sensitive and private information. Thus, directly releasing FMs trained on these data presents potential privacy issues because sensitive information encoded in FMs can possibly be recovered by adversaries. In order to share synthetic data generated by FMs without compromising the privacy of training data, privacy-preserving data generation methods have been proposed (Carranza et al., 2023; Yue et al., 2023; Mattern et al., 2022; Kurakin et al., 2023). Essentially, these methods first fine-tune a pretrained public FM using differential privacy (DP) on private domain data and then exploit the DP-tuned FM to generate new data, which can be distributed to downstream clients to augment their training data (illustrated in Figure 7(c)). These methods correspond to stage 1 of objective ① formulated in Eq.(1) but with an additional step of generating data using fine-tuned FMs.
@@ 5.2. FM adapted to DMs through representation-level knowledge transfer
Representation-level knowledge transfer (RLKT) refers to transferring and adapting knowledge of FMs to domain-specific models through passing representation-level knowledge, such as intermediate forward outputs and backward gradients, between FL participating parties. In literature, split learning (Liu et al., 2022) and representation transfer are two main categories of RLKT explored in the setting ① of FTL-FM. We summarize the two categories of RLKT methods in Table 4 and elaborate on these methods in this subsection.
@@
SAP (Shen et al., 2023) takes a similar setting as TextObfuscator, but it adopts a different training strategy: the client model is frozen while the server model is being fine-tuned using LoRA. To protect its data privacy, the client applies $d_{\chi}$-privacy (Chatzikokolakis et al., 2013) to obfuscate representation vectors before transmitting them to the server. Furthermore, SAP studied the impact of the FM’s split position on the privacy-preserving capacity and the model performance of the SAP framework. It empirically demonstrated that as the number of layers in the bottom model increases, the utility diminishes while the privacy-preserving capability strengthens, highlighting the existence of a trade-off between utility and privacy when partitioning the FM. SAP is a variant of the typical federated split learning we formulated in Eq.(9) as it conducts the FM adaptation through fine-tuning the model on the server side, as opposed to the client model.
@@ 5.4. Privacy and Backdoor Attacks
In the setting ①, the server is the potential adversary, who can launch privacy attacks with the intention of inferring clients’ private information. The server can also carry out backdoor attacks to compromise the utility of clients’ domain models.
@@ 6. Domain Knowledge for Augmenting Foundation Models
In this section, we overview FTL-FM methods that fall into setting ② (see Definition 3.1), the objective of which is to augment the FM with industry-level and domain-specific knowledge (e.g., healthcare, medicine, finance, banking, and law) transferred from clients.
@@
[…]
@@ 6.2. Domain knowledge transferred to FM through model-level knowledge transfer
To alleviate these limitations, Fan et al. (Fan et al., 2023; WeBank, [n. d.]) proposed a federated learning version of Offsite-Tuning named FedOST. In FedOST, the server first distributes the emulator and two adapters of its FM to all clients at the beginning of FL. Then, each client fine-tunes the two adapters on its private data with the assistance of the emulator. Next, all clients send their adapters to the server, which aggregates received adapters through secure aggregation (Bonawitz et al., 2016) and plugs the aggregated adapter into the FM. FedOST protects the privacy of clients’ private data and can build a fine-tuned FM with enhanced generalization capability. With similar motivations as FedOST, Kuang et al. integrated Offsite-Tuning into FederatedScope-LLM (Kuang et al., 2023b) and termed it FedOT. FedPEAT (Chua et al., 2023) generalized the Offsite-Tuning approach to federated learning with multiple communication rounds and proposed an adaptive control mechanism to facilitate the adoption of FedPEAT in a dynamic environment. CEFHRI (Khalid et al., 2023) adopted Offsite-Tuning to the human-robot interaction domain.